OpenStudy (anonymous):

Which of the following is an odd function? y = x^3 + 1 y = cosx y = sinx y = lnx y = e^-x

OpenStudy (anonymous):

y = x^3 + 1 y = sinx

OpenStudy (anonymous):

just look at their graphs

OpenStudy (anonymous):

use: www.wolframalpha.com

OpenStudy (anonymous):

@myko Well there can only be one answer. I'm going to assume it's sinx and not x^3+1, only due to the fact that it's +1 from the origin

OpenStudy (anonymous):

yes, you right . y = x^3 + is shifted, so it's not odd

OpenStudy (anonymous):

@myko Okay, thank you! :)